A balloon has a volume of 6.25 L at 453 kPa. What would the volume of this balloon be at 325 kPa


1
Expert's answer
2021-09-22T00:27:24-0400

Answer:

4.48 L

Explanation:

P1/V1 = P2/V2

453/6.25 = 325/ x

453x = 6.25*325

x = (6.25*325)/453

x = 4.48 L